a man borrow a sum of money and agree to pay it of by paying rs 3150 at the end of first year and ra 4400 at the end of second year .if the rate of simple interest is 5% per annum find the sum?
Answer by addingup(3677)

Simple interest: A= C(1+r). Now to your problem:
Solve for x:
1.05x-3150+0.05(1.05x-3150) = 4400 Let's simplify, eliminate the parenthesis:
0.05(1.05x-3150) = 0.0525x-157.5 OK, back to the equation but let's group like terms:
(1.05x+0.0525x)+(-3150-157.5) = 4400 Now add/subtract in this equation and you get:
1.1025x-3307.5 = 4400 Add 3307.05 on both sides:
1.1025x= 7707.05 Divide both sides by 1.1025:
x= 6,990.52 is the original amount the man borrowed.
-------------------------------------------------------------------
Check:
6990.52(1+0.05)= 6990.52*1.05= 7340.05 This is what the man owes at the end of the first year. At that point he pays 3150:
7340.05-3150= 4190.05 So we go into the second year:
4190.05*1.05= 4400 is the final payment on the debt. We have the correct answer (I get 43999.55 and it's because of the rounding of numbers)